# Surface hopping molecular dynamics simulation of ultrafast methyl iodide photodissociation mapped by Coulomb explosion imaging

### v1

@@ 3.1 Adiabatic potential energies of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$
The adiabatic PESs of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$ are constructed based on high-level ab initio electronic structure calculations and multi-dimensional cubic spline interpolations. Figure 1 shows the potential energies of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$ by stretching the C-I bond length $R$. Fifteen ionic potential curves with thresholds corresponding to different internal states of $\text{I}^{+}$ are shown in Fig. 1(a), and are also compared with the pure Coulomb interactions between $\text{CH}_{3}^{+}$ and $\text{I}^{+}$. The energies of the ionic states at the Franck-Condon point correspond to the double ionization potentials of $\text{CH}_{3}\text{I}$. These energies range from about 26 eV to 31 eV in our calculation, in good agreement with previous theoretical and experimental studies53, 54. The actual energies of all states are smaller than the pure Coulomb energy, owing to the attractive covalent interaction between $\text{CH}_{3}^{+}$ and $\text{I}^{+}$. This attractive interaction is particularly strong for small $R$. As the C-I distance increases, the potential becomes strongly repulsive and can be well approximated by pure Coulomb interactions for $R>4$.
Twelve neutral potential curves corresponding to the ground state and A-band excited states are shown in Fig. 1(b). At large C-I distance, these states dissociate to two atomic states of iodine, $\text{I}(^{2}\text{P}_{3/2})$ and $\text{I}^{*}(^{2}\text{P}_{1/2})$, with an energy splitting of 0.9 eV in our calculation. Using the Mulliken notation16, the ${}^{3}Q_{0}$, ${}^{1}Q_{1}$ and ${}^{3}Q_{1}$ states are dipole allowed transitions from the ground state. An UV pulse near 260 nm mostly excites the molecule to the ${}^{3}Q_{0}$ state20 through parallel transition, with a photoabsorption cross section much larger than the other states. It is noted that a non-adiabatic crossing between ${}^{3}Q_{0}$ and ${}^{1}Q_{1}$ states is present at $R=2.38$, which will induce population transfer to the ${}^{1}Q_{1}$ state during photodissociation.
@@
In previous CEI experiments, the measured delay-independent kinetic energy release (KER) show two major peaks centered around 4.4 eV and 5.2 eV for the doubly charged ion22, 27. Our simulation, as well as previous works22, 27, reveals that the KER peak near the 4.4 eV can be attributed to Coulomb explosions on the lowest ionic PES. [^4] The KER curves in Ref.22, 27 are mislabeled. The labels associated to the $\text{I}^{+}(^{3}\text{P}_{2})$ state and the $\text{I}^{+}(^{1}\text{D}_{2})$ state should be swapped. Therefore, the molecule is highly likely ionized to the lowest $\text{CH}_{3}\text{I}^{2+}$ ionic state corresponding to $\text{I}^{+}(^{3}\text{P}_{2})$ threshold after IR multi-photon absorption, and we simulate the Coulomb explosion part on this PES for most of our PD-CE trajectories. Figure 2(a) shows the lowest adiabatic PES of $\text{CH}_{3}\text{I}^{2+}$ as a function of the C-I distance $R$ and the angle $\alpha$ when the cation is in $C_{3v}$ symmetry. Contrary to the 1D potential curve shown in Fig.1(a) and in other literature22, 26, there is no local minimum on this PES. Consequently, our simulation finds that no metastable $\text{CH}_{3}\text{I}^{2+}$ cations can be formed for this ionic state, which is in contrast to the results using 1D wave packet propagation reported in Ref. 22. By further examination of all of the fifteen states in Fig. 1(a), only the state corresponding to the $\text{I}^{+}(^{1}\text{S}_{0})$ threshold has a shallow well near $R=2.25$, which may support metastable $\text{CH}_{3}\text{I}^{2+}$ cations as will be discussed with more details in Section 3.4.
Since all trajectories propagate on adiabatic PESs, we also show the PES of the 7th adiabatic state of $\text{CH}_{3}\text{I}$ in Fig. 2(b). This adiabatic state is separated into ${}^{3}Q_{0}$ and ${}^{1}Q_{1}$ states by a "seam" line, which indicates the geometries of the conical intersections. All PD-CE trajectories are initially prepared on this adiabatic PES near the Franck-Condon (FC) point, located at $R_{0}=2.16$ and $\alpha_{0}=107.6$°. As shown in Fig. 2(b), when the trajectories propagate close to the seam line, non-adiabatic transitions are likely to take place according to our Landau-Zener model and the trajectories may hop to the next adiabatic state.
@@ 3.2 Simulating the kinetic energy release
In a CEI experiment, the asymptotic momentum of each ionic fragment is measured via either velocity map imaging (VMI)14, 22, 23, 24, 25, 26 or COLd Target Recoil-Ion Momentum Spectrometer (COLTRIMS)9, 11, 12, 13, 15, 27. In the latter case, the measured kinetic energy release (KER) is the sum of the kinetic energies of $\text{I}^{+}$ and $\text{CH}_{3}^{+}$ fragments detected in coincidence. In our simulation, this corresponds to the translational kinetic energy of the relative motion between I($\text{I}^{+}$) and $\text{CH}_{3}$($\text{CH}_{3}^{+}$) fragments, which can be written as
$KER=T_{trans}=\frac{1}{2}\mu\dot{\textbf{R}}_{I-CH_{3}}^{2},$ (5)
where $\mu$ is the reduced mass between I($\text{I}^{+}$) and $\text{CH}_{3}$($\text{CH}_{3}^{+}$) fragments, $\textbf{R}_{I-CH_{3}}$ is the displacement vector between I($\text{I}^{+}$) and the center of mass of $\text{CH}_{3}$($\text{CH}_{3}^{+}$), and $\dot{\textbf{R}}_{I-CH_{3}}=d\textbf{R}_{I-CH_{3}}/dt$ is the velocity of the relative motion.
@@
$KER=KER_{PD}+KER_{CE},$ (6)
where $KER_{PD}$ and $KER_{CE}$ are the kinetic energies gained from photodissociation and Coulomb explosion, respectively.
@@
When a one-dimensional (1D) model is applied to the PD-CE process, that is, considering the C-I distance $R$ as the only reaction coordinate and freezing all other degrees of freedom at the equilibrium geometry, $KER_{PD}$ and $KER_{CE}$ are approximated to be
$\begin{split}KER_{PD}(t)\approx V_{\text{CH}_{3}\text{I}}(R_{0})-V_{\text{CH}_{3}\text{I}}(R(t)),\\ KER_{CE}(t)\approx V_{\text{CH}_{3}\text{I}^{2+}}(R(t))-V_{\text{CH}_{3}\text{I}^{2+}}(\infty),\\ t=\int_{R_{0}}^{R}\frac{d\tilde{R}}{\sqrt{2(V_{\text{CH}_{3}\text{I}}(R_{0})-V_{\text{CH}_{3}\text{I}}(\tilde{R}))/\mu}},\end{split}$ (7)
where $t$ is the pump-probe delay. This 1D model is used for interpreting the experimental data in Ref.27. At large pump-probe delays, the C-I distance $R$ is large and the photodissociation is complete. In that case, $KER_{PD}\approx V_{\text{CH}_{3}\text{I}}(R_{0})-V_{\text{CH}_{3}\text{I}}^{th}$ is approximately a constant value and $KER_{CE}\approx 1/R(t)$ is approximately the pure Coulomb energy.
@@ 3.3 Coulomb explosion in the asymptotic region
Fig.  4: Kinetic energy release (KER) at different pump-probe delays ranging from 60 fs to 500 fs. (a) Signals from our simulation of about 20,000 PD-CE trajectories at T=200 K. The KER equals to the final translational kinetic energy of the relative motion between $\text{CH}_{3}^{+}$ and $\text{I}^{+}$ fragments. (b) Signals adapted from the experiment in Ref.27. The KER is the sum of the kinetic energies of $\text{CH}_{3}^{+}$ and $\text{I}^{+}$ fragments measured in a coincident manner for a UV pump intensity of $1\times 10^{13}\text{W}/\text{cm}^{2}$. The yellow and black curves are predictions using the 1D model in Eq.(7). Signals in both (a) and (b) are counted using a bin size $h_{t}=20$ fs and $h_{E}=0.01$ eV.
@@
About 20,000 PD-CE trajectories are calculated at T=200 K with the pump-probe delay ranging from 50 fs to 500 fs and with a step of $\Delta t=1$ fs. Since previous studies have concluded $\text{CH}_{3}\text{I}$ will be highly likely ionized to the lowest ionic states after IR multi-photon absorption22, 23, the Coulomb explosion part is simulated on the lowest PES corresponding to the $\text{I}^{+}(^{3}\text{P}_{2})$ threshold. The simulated delay-dependent KER signals are shown in Fig. 4(a), along with the relation between KER and the pump-probe delay predicted by the 1D model in Eq.(7). The signals form two major bands, corresponding to the I and $\text{I}^{*}$ reaction channels. The lower band ($\text{I}^{*}$ channel) is significantly stronger than the upper band (I channel), indicating that a large fraction of trajectories undergo non-adiabatic transitions. The branching ratio $[\text{I}^{*}]/([\text{I}^{*}]+[\text{I}])$ is 0.82 in our simulation. This value is in overall agreement with simulations using the FSSH algorithm32, 31, wave packet evolutions35 and experiments56, 57, 58, 59, which range from 0.65 to 0.85.
@@
The two bands show clear correlation between the KER and the pump-probe delay, which agree with the prediction of the 1D model. The energy of the upper band is slightly lower than that of the 1D model, because more potential energy is converted to the vibrational and rotational motion of $\text{CH}_{3}$ fragment in $\text{I}^{*}$ reaction channel. In fact, over ninety percent of the potential energy is redistributed to the translational energy of $\text{CH}_{3}$ and I fragments during photodissociation, while a small fraction is redistributed to the vibrational and rotational motion of $\text{CH}_{3}$. Previous studies reveal that higher vibrational and rotational quantum states of $\text{CH}_{3}$ are excited in $\text{I}^{*}$ dissociation channel31, 33.
@@
The simulated delay-dependent KER signals are also in good agreement with the experiment, as shown in Fig.4(b), although the experiment shows slightly lower KER than the 1D model. We only compare our simulation with the experiment for KER below 4 eV because higher KER region in the experiment is dominated by delay-independent signals due to multi-photon ionization of ground-state $\text{CH}_{3}\text{I}$ by the probe pulse, which is beyond the focus of this work. A weak contribution of lower KER in the experiment is attributed to multi-photon Rydberg state excitation by the pump pulse, which is not accounted for in our simulation.
@@
Fig.  5: Experimental kinetic energy release (KER) signals (dashed black, adapted from Ref.27) compared with simulated $KER_{PD}$ signals (solid black). The experimental signals are integrated over the pump-probe delay from 2.5 ps to 3ps. The intensity of the UV pump is $1\times 10^{13}\text{W}/\text{cm}^{2}$. The simulated $KER_{PD}$ include only the kinetic energy release during photodissociation, and are integrated over the pump-probe delay from 50 fs to 100 fs. The blue and red curves with shadows are separate signals corresponding to $\text{I}^{*}$ and I dissociation channels respectively in our simulation, scaled by a factor of 1/2. The vertical lines indicate the potential energy differences between ${}^{3}\text{Q}_{0}$ state at equilibrium geometry and the thresholds of $\text{I}^{*}$(red) and I(blue) channels, with the relative heights matching the $[\text{I}^{*}]/([\text{I}^{*}]+[\text{I}])$ branching ratio in our simulation. Signals are normalized by the total number of events between 1 eV and 3 eV.
@@
When the pump-probe delay is sufficiently large such that the Coulomb energy is negligible ($KER_{CE}$<0.05 eV when t>2.5 ps), the measured KER is entirely contributed from $KER_{PD}$ during photodissociation, which has already saturated at t=50 fs. Therefore, it is convenient to compare the simulated $KER_{PD}$ at t>50 fs with the measured KER at t>2.5 ps, as shown in Fig. 5. The overall shape and peak positions of the simulated $\text{KER}_{PD}$ signals shows good agreement with the experimentally measured KER. The relative height between the two peaks in our simulation is slightly different from the experiment, which reflects the branching ratio in our simulation is higher than the experiment. This can be attributed to the limitation of the present surface hopping model. Another reason is that we neglect the small fraction of dipole transitions to the ${}^{1}\text{Q}_{1}$ state that mostly leads to dissociations in I channel.

### v2

@@ 3.1 Adiabatic potential energies of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$
The adiabatic PESs of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$ are constructed based on high-level ab initio electronic structure calculations and multivariate cubic spline interpolations. Figure 1 shows the potential energies of $\text{CH}_{3}\text{I}$ and $\text{CH}_{3}\text{I}^{2+}$ by stretching the C-I bond length $R$. Fifteen ionic potential curves with thresholds corresponding to different internal states of $\text{I}^{+}$ are shown in Fig. 1(a), and are also compared with the pure Coulomb interactions between $\text{CH}_{3}^{+}$ and $\text{I}^{+}$. The energies of the ionic states at the Franck-Condon point correspond to the double ionization potentials of $\text{CH}_{3}\text{I}$. These energies range from about 26 eV to 31 eV in our calculation, in good agreement with previous theoretical and experimental studies60, 61. The actual energies of all states are smaller than the pure Coulomb energies, owing to the attractive covalent interaction between $\text{CH}_{3}^{+}$ and $\text{I}^{+}$. This attractive interaction is particularly strong for small $R$. As the C-I distance increases, the potential becomes strongly repulsive and can be well approximated by pure Coulomb interactions for $R>4$.
Twelve neutral potential curves corresponding to the ground state and A-band excited states are shown in Fig. 1(b). At large C-I distance, these states dissociate to two atomic states of iodine, $\text{I}(^{2}\text{P}_{3/2})$ and $\text{I}^{*}(^{2}\text{P}_{1/2})$, with an energy splitting of 0.9 eV in our calculation. Using the Mulliken notation19, the ${}^{3}\text{Q}_{0}$, ${}^{1}\text{Q}_{1}$ and ${}^{3}\text{Q}_{1}$ diabatic states are dipole allowed transitions from the ground state. An UV pulse near 260 nm mostly excites the molecule to the ${}^{3}\text{Q}_{0}$ state23 through parallel transition. It is noted that a non-adiabatic crossing between ${}^{3}\text{Q}_{0}$ and ${}^{1}\text{Q}_{1}$ states is present at $R=2.38$, which will induce population transfer to the ${}^{1}\text{Q}_{1}$ state during photodissociation.
@@
In previous CEI experiments, the measured delay-independent kinetic energy release (KER) show two major peaks centered around 4.4 eV and 5.2 eV for the doubly charged ion25, 30. Our simulation, as well as previous works25, 30, reveals that the KER peak near the 4.4 eV can be attributed to Coulomb explosions on the lowest ionic PES. [^4] The KER curves in Ref.25, 30 are mislabeled. The labels associated to the $\text{I}^{+}(^{3}\text{P}_{2})$ state and the $\text{I}^{+}(^{1}\text{D}_{2})$ state should be swapped. Therefore, the molecule is highly likely ionized to the lowest $\text{CH}_{3}\text{I}^{2+}$ ionic state corresponding to $\text{I}^{+}(^{3}\text{P}_{2})$ threshold after IR multi-photon absorption, and we simulate the Coulomb explosion part on this PES for most of our PD-CE trajectories. Moreover, conical intersections are absent on this PES, justifying our simplification of not considering non-adiabatic effects for the Coulomb explosion part. Figure 2(a) shows the lowest adiabatic PES of $\text{CH}_{3}\text{I}^{2+}$ as a function of the C-I distance $R$ and the angle $\alpha$ when the cation is in $C_{3v}$ symmetry. Contrary to the 1D potential curve shown in Fig.1(a) and in other literature25, 29, there is no local minimum on this PES. Consequently, our simulation finds that no metastable $\text{CH}_{3}\text{I}^{2+}$ cations can be formed for this ionic state, which is in contrast to the results using 1D wave packet propagation reported in Ref. 25. By further examination of all of the fifteen states in Fig. 1(a), only the state corresponding to the $\text{I}^{+}(^{1}\text{S}_{0})$ threshold has a shallow well near $R=2.25$, which may support metastable $\text{CH}_{3}\text{I}^{2+}$ cations as will be discussed with more details in Section 3.4.
Since all trajectories propagate on adiabatic PESs, we also show the PES of the 7th adiabatic state of $\text{CH}_{3}\text{I}$ in Fig. 2(b). This adiabatic state is separated into ${}^{3}\text{Q}_{0}$ and ${}^{1}\text{Q}_{1}$ diabatic states by a "seam" line, which indicates the geometries of the conical intersections. All PD-CE trajectories are initially prepared on this adiabatic PES near the Franck-Condon (FC) point, located at $R_{0}=2.16$ and $\alpha_{0}=107.6$°. As shown in Fig. 2(b), when the trajectories propagate close to the seam line, non-adiabatic transitions are likely to take place according to the Landau-Zener model and the trajectories may hop to a different adiabatic state.
@@ 3.2 Simulating the kinetic energy release
In a CEI experiment, the asymptotic momentum of each ionic fragment is measured via either velocity map imaging (VMI)15, 25, 26, 27, 28, 29 or COLd Target Recoil-Ion Momentum Spectrometer (COLTRIMS)9, 11, 12, 13, 16, 30. In the latter case, the measured kinetic energy release (KER) is the sum of the kinetic energies of $\text{I}^{+}$ and $\text{CH}_{3}^{+}$ fragments detected in coincidence. In our simulation, this corresponds to the translational kinetic energy of the relative motion between I($\text{I}^{+}$) and $\text{CH}_{3}$($\text{CH}_{3}^{+}$) fragments, which can be written as
$\mathrm{KER}=\mathrm{T_{trans}}=\frac{1}{2}\mu\dot{\textbf{R}}_{\mathrm{I-CH_{3}}}^{2},$ (6)
where $\mu$ is the reduced mass between I($\text{I}^{+}$) and $\text{CH}_{3}$($\text{CH}_{3}^{+}$) fragments, $\textbf{R}_{\mathrm{I-CH_{3}}}$ is the displacement vector between I($\text{I}^{+}$) and the center of mass of $\text{CH}_{3}$($\text{CH}_{3}^{+}$), and $\dot{\textbf{R}}_{\mathrm{I-CH_{3}}}=d\textbf{R}_{\mathrm{I-CH_{3}}}/dt$ is the velocity of the relative motion.
@@
$\mathrm{KER}=\mathrm{KER_{PD}}+\mathrm{KER_{CE}},$ (7)
where $\mathrm{KER_{PD}}$ and $\mathrm{KER_{CE}}$ are the kinetic energies gained from photodissociation and Coulomb explosion, respectively.
@@
When a one-dimensional (1D) model is applied to the PD-CE process, that is, considering the C-I distance $R$ as the only reaction coordinate and freezing all other degrees of freedom at the equilibrium geometry, $\mathrm{KER_{PD}}$ and $\mathrm{KER_{CE}}$ are approximated to be
$\begin{split}\mathrm{KER_{PD}}(t)\approx V_{\text{CH}_{3}\text{I}}(R_{0})-V_{\text{CH}_{3}\text{I}}(R(t)),\\ \mathrm{KER_{CE}}(t)\approx V_{\text{CH}_{3}\text{I}^{2+}}(R(t))-V_{\text{CH}_{3}\text{I}^{2+}}(\infty),\\ t=\int_{R_{0}}^{R}\frac{d\tilde{R}}{\sqrt{2(V_{\text{CH}_{3}\text{I}}(R_{0})-V_{\text{CH}_{3}\text{I}}(\tilde{R}))/\mu}},\end{split}$ (8)
where $t$ is the pump-probe delay. This 1D model is used for interpreting the experimental data in Ref.30. At large pump-probe delays, the C-I distance $R$ is large and the photodissociation is complete. In that case, $\mathrm{KER_{PD}}\approx V_{\text{CH}_{3}\text{I}}(R_{0})-V_{\text{CH}_{3}\text{I}}^{th}$ is approximately a constant value and $\mathrm{KER_{CE}}\approx 1/R(t)$ is approximately the pure Coulomb energy.
@@ 3.3 Coulomb explosion in the asymptotic region
Fig.  4: Kinetic energy release (KER) at different pump-probe delays ranging from 60 fs to 500 fs. (a) Signals from our simulation of about 20,000 PD-CE trajectories at $T=200$ K. The KER equals to the final translational kinetic energy of the relative motion between $\text{CH}_{3}^{+}$ and $\text{I}^{+}$ fragments. (b) Signals adapted from the experiment in Ref.30. The KER is the sum of the kinetic energies of $\text{CH}_{3}^{+}$ and $\text{I}^{+}$ fragments measured in a coincident manner for a UV pump intensity of $1\times 10^{13}\text{W}/\text{cm}^{2}$. The yellow and black curves are predictions using the 1D model in Eq.(8). Signals in both (a) and (b) are counted using a bin size $h_{t}=20$ fs and $h_{E}=0.01$ eV.
@@
About 20,000 PD-CE trajectories are calculated at $T=200$ K with the pump-probe delay ranging from 50 fs to 500 fs and with a step of $\Delta t=1$ fs. Since previous studies have concluded that $\text{CH}_{3}\text{I}$ will be highly likely ionized to the lowest ionic states after IR multi-photon absorption25, 26, the Coulomb explosion part is simulated on the lowest PES corresponding to the $\text{I}^{+}(^{3}\text{P}_{2})$ threshold. The resulting simulated delay-dependent KER signals are shown in Fig. 4(a), along with the relation between KER and the pump-probe delay predicted by the 1D model in Eq.(8). The signals form two major bands, corresponding to the I and $\text{I}^{*}$ reaction channels. The lower band ($\text{I}^{*}$ channel) is significantly stronger than the upper band (I channel), indicating that a large fraction of trajectories undergo non-adiabatic transitions. The branching ratio $[\text{I}^{*}]/([\text{I}^{*}]+[\text{I}])$ is 0.82 in our simulation. This value is in overall agreement with simulations using the FSSH algorithm35, 34, wave packet evolutions38 and experiments63, 64, 65, 66, which range from 0.65 to 0.85.
@@
The two bands show clear correlation between the KER and the pump-probe delay, which agree with the prediction of the 1D model. The energy of the upper band is slightly lower than that of the 1D model, because more potential energy is converted to the vibrational and rotational motion of $\text{CH}_{3}$ fragment in the $\text{I}^{*}$ reaction channel. In fact, over ninety percent of the potential energy is redistributed to the translational energy of $\text{CH}_{3}$ and I fragments during photodissociation, while a small fraction is redistributed to the vibrational and rotational motion of $\text{CH}_{3}$. Previous studies reveal that higher vibrational and rotational quantum states of $\text{CH}_{3}$ are excited in $\text{I}^{*}$ dissociation channel34, 36.
@@
The simulated delay-dependent KER signals are also in good agreement with the experiment, as shown in Fig.4(b), although the experiment shows slightly lower KER than the 1D model. We only compare our simulation with the experiment for KER below 4 eV because higher KER region in the experiment is dominated by delay-independent signals due to multi-photon ionization of ground-state $\text{CH}_{3}\text{I}$ by the probe pulse, which is beyond the focus of this work. A weak contribution at lower KER in the experiment is attributed to multi-photon Rydberg state excitation by the pump pulse, which is not accounted for in our simulation. Unlike the simulated signals that are evenly distributed in each time interval, more photodissociation signals are detected in the experiment at larger pump-probe delays, which leads to the intensity enhancement of the lower KER band as a function of time. The most likely reason for this is that the strong-field dissociative double ionization rate increases as a function of the internuclear distance $R$. Such effects are well known, e.g. due to a varying ionization potential as a function of internuclear separation or due to resonance-like effects dubbed “enhanced ionization”67, 68. Since we do not model the strong-field ionization process, such effects are absent in our simulation, and the ion pair yield in the simulation is, by definition, constant.
@@
Fig.  5: Experimental kinetic energy release (KER) signals (solid black, adapted from Ref.30) compared with simulated $\mathrm{KER_{PD}}$ signals (dashed black). The experimental signals are integrated over the pump-probe delay from 2.5 ps to 3 ps for a UV pump intensity of $1\times 10^{13}\text{W}/\text{cm}^{2}$. The simulated $\mathrm{KER_{PD}}$ include only the kinetic energy release during photodissociation, and are integrated over the propagation time from 50 fs to 100 fs. The blue and red curves are separate signals corresponding to I and $\text{I}^{*}$ dissociation channels respectively in our simulation. The vertical lines indicate the potential energy differences between ${}^{3}\text{Q}_{0}$ state at the equilibrium geometry and the thresholds of $\text{I}^{*}$(red) and I(blue) channels, with the relative heights matching the $[\text{I}^{*}]/([\text{I}^{*}]+[\text{I}])$ branching ratio in our simulation. Signals are normalized by the total number of events between 1 eV and 3 eV.
When the pump-probe delay is sufficiently large such that the Coulomb energy is negligible ($\mathrm{KER_{CE}}(t>2.5\text{ ps})\approx 0$), the measured KER is entirely due to the $\mathrm{KER_{PD}}$ during photodissociation. The $\mathrm{KER_{PD}}$ has already saturated at a propagation time $t=50$ fs ($\mathrm{KER_{PD}}(t>50\text{ fs})\approx\text{Constant}$), as demonstrated in Fig. 3. Therefore, we can write the KER in this regime as follows:
$\begin{split}\mathrm{KER}(t>2.5\text{ ps})&=\mathrm{KER_{CE}}(t>2.5\text{ ps})+\mathrm{KER_{PD}}(t>2.5\text{ ps})\\ &\approx\mathrm{KER_{PD}}(t>50\text{ fs})\\ &\approx\text{Constant}\end{split}$ (9)
@@
Therefore, it is convenient to compare the simulated $\mathrm{KER_{PD}}$(t>50 fs) with the measured KER(t>2.5 ps), as shown in Fig. 5. The overall shape and peak positions of the simulated $\mathrm{KER_{PD}}$ signals shows good agreement with the experimentally measured KER. The relative height between the two peaks in our simulation is slightly different from the experiment, which reflects that the branching ratio in our simulation is higher than in the experiment. This can be attributed to the limitation of the present surface hopping model. Another reason is that we neglect the small fraction of dipole transitions to the ${}^{1}\text{Q}_{1}$ state that mostly leads to dissociations into the I channel.